OpenStudy (anonymous):

Elimination!!! Solving systems of equations... Can anyone help me figure out these, using elimination? 3x-6y=3 7x-5y=-1 Thank yous!

OpenStudy (anonymous):

Multiply the first equation by -(7/3) and get -(7/3)*3x + (7/3)*6y = -(7/3)*3 -7x + 14y = -7 Now add first and second equations and find -7x + 7x - 5y + 14y = -7 -1 9y = -8 y = -8/9 Substitute back and find x

OpenStudy (anonymous):

It's a bit simpler to reduce the 3x-6y = 3 --> x -2y = 1 first then multiply by -7

OpenStudy (anonymous):

okay... thanks. Both of ya! ;P